# Ellipsoidal Density-Equalizing Map for Genus-0 Closed Surfaces

### v1

@@ 14      Update the mapping $f_{n+1}=f_{n}+\delta tdE^{n}_{1}$;
21             Apply the shape update method (Algorithm 2) with the radius step sizes $0.9^{m}\delta b$ and $0.9^{m}\delta c$ and obtain the updated elliptic radii $(\tilde{a},\tilde{b},\tilde{c})$;
22
23            Set $(a,b,c)=(\tilde{a},\tilde{b},\tilde{c})$;
24
@@ 26      Update $n=n+1$;
27
28until $|\frac{E^{n+1}-E^{n}}{E^{n}}|<\epsilon$  or  $n\geq n_{\max}$;
29The resulting bijective ellipsoidal density-equalizing quasi-conformal map is $f:\mathcal{M}\rightarrow\mathcal{E}_{a,b,c}$, where $(a,b,c)$ are the optimal elliptic radii;
@@ 4.1 Ellipsoidal density-equalizing map
Figure 4: Ellipsoidal density-equalizing maps of ellipsoidal surfaces. Each row shows one example. (a) An example with discontinuous input density. (b) An example with continuous input density. Left to right: The initial ellipsoidal surface color-coded with the initial density, the final EDEM result color-coded with the initial density, the histogram of the initial density, and the histogram of the final density.
@@
Figure 5: Additional examples of ellipsoidal density-equalizing maps of ellipsoidal surfaces. For each example, the left figure shows the input surface color-coded with the initial density, and the right figure shows the EDEM result color-coded with the initial density. (a) An ellipsoid with non-uniformly distributed mesh elements and a prescribed discontinuous density. (b) An elongated ellipsoid with a prescribed continuous density. (c) An ellipsoid with non-uniformly distributed mesh elements and a complex prescribed density. (d) A sphere with a prescribed continuous density.
@@
Surface # Faces Time (s) $\mathbf{(a,b,c)}$ $\text{Var}(\widetilde{\rho}_{1})$ $\text{Var}(\widetilde{\rho}_{2})$ Overlaps
@@
Next, we consider computing the ellipsoidal density-equalizing maps for more general genus-0 closed surfaces. In Fig. 6, we set the population as the face area of the original mesh and apply our EDEM algorithm to achieve ellipsoidal area-preserving parameterizations. For the initial parameterization step, we map each surface conformally onto the ellipsoid using the FECM method [30]. Then, we apply the EDEM method to compute the ellipsoidal density-equalizing parameterizations. To assess the area-preserving property of the ellipsoidal parameterizations, we consider the area distortion for each triangular face $T$ as follows:
@@
Here $f$ denotes the ellipsoidal parameterization and $\mathcal{F}$ is the set of all triangular faces. The normalization factors in the numerator and denominator ensure that a perfectly area-preserving mapping would result in $d_{\text{area}}\equiv 0$. In Fig. 6, we present various genus-0 surface models, the initial ellipsoidal parameterizations, the ellipsoidal parameterizations obtained by our EDEM method, and the initial and final area distortions. It can be observed that the EDEM method performs very well for genus-0 closed surfaces with complex structures. Specifically, by visually comparing the triangle element distributions in the initial ellipsoidal parameterizations and the EDEM parameterization results, it is easy to see that the EDEM method achieves a more uniform distribution on the prescribed ellipsoid. By comparing the initial and final area distortion histograms, we can also see that the area distortion is significantly reduced by our EDEM method. For a more quantitative analysis, Table 2 provides detailed statistics on the performance of our EDEM algorithm for ellipsoidal area-preserving parameterization. We can see that the EDEM method can efficiently reduce the area distortion by over 95% on average when compared to the initial parameterization, and the bijectivity of the parameterization is well-preserved in all examples. Altogether, the experiments show that our EDEM method is capable of computing bijective ellipsoidal area-preserving parameterizations for a large variety of genus-0 closed surfaces.
Figure 6: Ellipsoidal area-preserving parameterization of genus-0 closed surfaces obtained by the EDEM method. Each row shows one example. (a) The David model. (b) The Buddha model. (c) The hippocampus model. (d) The twisted ball model. Left to right: The input surface mesh, the initial ellipsoidal conformal parameterization obtained by the FECM method [30], the EDEM result, the histogram of the logged area ratio $d_{\text{area}}$ of the initial ellipsoidal parameterization, and the histogram of the logged area ratio $d_{\text{area}}$ of the final ellipsoidal parameterization.
Table 2: The performance of our EDEM algorithm for ellipsoidal area-preserving parameterization. For each surface, we record the number of triangle elements, the computational time, the elliptic radii $(a,b,c)$ of the target ellipsoidal domain, mean and standard deviation of the initial area distortion $|d_{\text{area}}(f_{0})|$ and the final area distortion $|d_{\text{area}}(f)|$, and the number of overlaps.
Surface # Faces Time (s) $\mathbf{(a,b,c)}$ $|d_{\text{area}}(f_{0})|$ $|d_{\text{area}}(f)|$ Overlaps
@@
Table 3: The results obtained by our EDEM algorithm with different elliptic radii. Here, for different choices of the elliptic radii $(a,b,c)$ of the target ellipsoid, we record the mean and standard deviation of the final area distortion $|d_{\text{area}}(f)|$, the mean value of the norm of the Beltrami coefficient $|\mu|$, and the number of overlaps.
$\mathbf{a}$ $\mathbf{b}$ $\mathbf{c}$ $\text{Mean}(|d_{\text{area}}(f)|)$ $\text{SD}(|d_{\text{area}}(f)|)$ Mean$(|\mu|)$ Overlaps
@@
It is natural to ask whether the performance of our EDEM method is affected by the shape of the target ellipsoid. Here, we map the Hippocampus model to various target ellipsoids with different elliptic radii using our EDEM method and analyze the area and angle distortions of the resulting mappings. In Table 3, we report the mean and variance of the area distortion, as well as the mean value of the norm of the Beltrami coefficient $|\mu|$. It can be observed that for all combinations of the elliptic radii, all mapping results are folding-free and possess low area distortion. This shows that our EDEM method is capable of computing area-preserving parameterizations of genus-0 closed surfaces onto different prescribed ellipsoids. However, if we also take the angle distortion into consideration, then some differences can be observed. For instance, if we increase the value of $c$, we can see that the angle distortion is further reduced. This experiment suggests that the overall ellipsoidal shape plays an important role and motivates the need for the proposed EDEQ method, which optimizes the ellipsoidal geometry and achieves ellipsoidal density-equalizing quasi-conformal maps.
Figure 7: Ellipsoidal density-equalizing quasi-conformal maps of ellipsoidal surfaces. Each row shows one example. (a) An example with discontinuous input density. (b) An example with continuous input density. Left to right: The initial ellipsoidal surface color-coded with the initial density, the final EDEQ result color-coded with the initial density, the histogram of the initial density, and the histogram of the final density.
@@ 4.2 Ellipsoidal density-equalizing quasi-conformal map
Next, we test our EDEQ algorithm for computing the ellipsoidal density-equalizing quasi-conformal mappings. In Fig. 7(a), we define different populations at different regions on the ellipsoid to obtain a discontinuous initial density. Note that the maximum density is three times larger than the minimum one. It can be observed in the EDEQ mapping result that the shape of the ellipsoid is changed under the mapping process, while the higher-density domain expands and the lower-density domain contracts. Comparing the initial and final density histograms, it can be observed that the density is highly equalized. Fig. 7(b) shows another example with a continuous initial density defined on an ellipsoid. Under the EDEQ algorithm, the ellipsoidal shape is stretched along the z-direction and becomes more elongated, and different regions are expanded or shrunk according to the given density. From the density histograms, it can again be observed that the mapping result is highly density-equalizing.
@@
Figure 8: Ellipsoidal density-equalizing quasi-conformal maps of ellipsoidal surfaces. For each example, the left figure shows the input surface color-coded with the initial density, and the right figure shows the EDEQ result color-coded with the initial density. (a) An ellipsoid with non-uniformly distributed mesh elements and a prescribed discontinuous density. (b) An elongated ellipsoid with a prescribed continuous density. (c) An ellipsoid with non-uniformly distributed mesh elements and a complex prescribed density. (d) A sphere with a prescribed continuous density.
@@
In Fig. 8, we further apply the EDEQ method to the same set of ellipsoidal examples in Fig. 5. It can be observed that our EDEQ method produces ellipsoidal density-equalizing mapping results with both the ellipsoidal geometry and the vertex positions optimized. For a more quantitative analysis, we record the initial and final elliptic radii, the variance of the initial and final densities, the mean value of the Beltrami coefficient $\mu$, and the number of overlaps for all the above-mentioned examples in Table 4. From the table, we can see that the ellipsoidal shapes are effectively optimized under the EDEQ method. Furthermore, in addition to possessing a highly density-equalizing effect as reflected in the variance of the final density, the mapping results also achieve a low conformal distortion. Specifically, note that the norm of the Beltrami coefficient $|\mu|$ normally ranges from 0 to 1, where $|\mu|=0$ indicates that the mapping is conformal. As the value of $|\mu|$ is small for all examples, we can see that the EDEQ method not only achieves density equalization but also reduces the conformal distortion of the ellipsoidal mappings.
@@
Next, we consider using the EDEQ method to compute the ellipsoidal area-preserving parameterizations for genus-0 closed surfaces. For each surface in Fig. 9, we first conformally map it onto an ellipsoid with some arbitrary elliptic radii using the FECM method [30]. Then, we apply the EDEQ algorithm with the population set as the original face area to obtain the ellipsoidal area-preserving parameterizations. It can be observed that the ellipsoidal shapes are changed significantly under the EDEQ method. Comparing the initial and final area distortion histograms, we can see that the mapping results are highly area-preserving.
Figure 9: Ellipsoidal area-preserving quasi-conformal parameterization of genus-0 closed surfaces. Each row shows one example. (a) The Duck model. (b) The Hippocampus model. (c) The Lion-Vase model. Left to right: The input surface mesh, the initial ellipsoidal conformal parameterization, the final EDEQ result, the histogram of the logged area ratio $d_{\text{area}}$ of the initial ellipsoidal parameterization, and the histogram of the logged area ratio $d_{\text{area}}$ of the final ellipsoidal parameterization.
@@
As the choice of the initial elliptic radii is arbitrary, we can even start with a unit sphere and run the EDEQ method to eventually obtain an optimal ellipsoidal parameterization. In Fig. 10, we consider several genus-0 closed surface models and compute an initial spherical conformal parameterization. Then, we apply the proposed EDEQ algorithm to compute an optimal ellipsoidal area-preserving parameterization. It can be observed that throughout the EDEQ iterative process, both the ellipsoidal shapes and the vertex positions on the ellipsoids are changed gradually. For a more quantitative analysis, Table 5 summarizes the performance of our EDEQ algorithm in terms of the final radii, the initial and final area distortions, the conformal distortion, and the number of overlaps. Comparing the initial and final area distortions, it can be observed that the area distortion is significantly reduced in all examples. The mean of the norm of the Beltrami coefficient $|\mu|$ is also small, indicating that the conformal distortion is low. Besides, the number of overlaps is 0 for all examples, which shows that our mapping results are all folding-free. In Table 6, We further compare the performance of the EDEQ method and the EDEM method in terms of the area and conformal distortions, where both methods use the same initial elliptic radii. It can be observed that the EDEQ method achieves a significantly lower conformal distortion when compared to the EDEM method, while the area distortion remains comparable. Altogether, the above experiments demonstrate the effectiveness of our proposed EDEQ method for ellipsoidal density-equalizing quasi-conformal maps.
Table 5: The performance of our EDEQ algorithm for ellipsoidal area-preserving parameterization. For each surface, we start with an initial spherical conformal parameterization and then apply the EDEQ algorithm. The number of triangle elements, the final elliptic radii $(a,b,c)$, the mean and standard deviation of the initial area distortion $|d_{\text{area}}(f_{0})|$ and the final area distortion $|d_{\text{area}}(f)|$, the mean value of the norm of the Beltrami coefficient $|\mu|$, and the number of overlaps are recorded.
@@
Mean of $|d_{\text{area}}(f)|$ SD of $|d_{\text{area}}(f)|$ Mean$(|\mu|)$ Overlaps Mean of $|d_{\text{area}}(f)|$ SD of $|d_{\text{area}}(f)|$ Mean$(|\mu|)$ Overlaps

### v2

@@ 15  Update the mapping $f_{n+1}=f_{n}+\delta tdE^{n}_{1}$;
[…]
23
[…]
@@ 27  Update $n=n+1$;
28
29until $|\frac{E^{n+1}-E^{n}}{E^{n}}|<\epsilon$ or $n\geq n_{\max}$;
30The resulting bijective ellipsoidal density-equalizing quasi-conformal map is $f:\mathcal{M}\rightarrow\mathcal{E}_{a,b,c}$, where $(a,b,c)$ are the optimal elliptic radii;
@@ 4.1 Ellipsoidal density-equalizing map
Figure 4: Ellipsoidal density-equalizing maps of ellipsoidal surfaces. Each row shows one example. (a) An example with discontinuous input density. (b) An example with continuous input density. Left to right: The initial ellipsoidal surface color-coded with the initial density, the final EDEM result color-coded with the initial density, the histogram of the initial density, and the histogram of the final density. The elliptic radii are $(a,b,c)=(1,2,4)$ for both examples.
@@
Figure 5: Additional examples of ellipsoidal density-equalizing maps of ellipsoidal surfaces. For each example, the left figure shows the input surface color-coded with the initial density, and the right figure shows the EDEM result color-coded with the initial density. (a) An ellipsoid with non-uniformly distributed mesh elements and a prescribed discontinuous density. Here, the elliptic radii are $(a,b,c)=(1,1,1.5)$. (b) An elongated ellipsoid with a prescribed continuous density. Here, the elliptic radii are $(a,b,c)=(1,0.6,2)$. (c) An ellipsoid with non-uniformly distributed mesh elements and a complex prescribed density. Here the radii are $(a,b,c)=(1,1,1.5)$. (d) A sphere with a prescribed continuous density. Here, the elliptic radii are $(a,b,c)=(1,1,1)$.
@@
Surface # Faces Time (s) $\mathbf{(a,b,c)}$ $\text{Var}(\widetilde{\rho}_{1})$ $\text{Var}(\widetilde{\rho}_{2})$ # Overlaps
@@
Next, we consider computing the ellipsoidal density-equalizing maps for more general genus-0 closed surfaces. In Fig. 6, we set the population as the face area of the original mesh and apply our EDEM algorithm to achieve ellipsoidal area-preserving parameterizations. For the initial parameterization step, we map each surface conformally onto the ellipsoid using the FECM method [30]. Then, we apply the EDEM method to compute the ellipsoidal density-equalizing parameterizations. To assess the area-preserving property of the ellipsoidal parameterizations, we consider the log-area error for each triangular face $T$ as follows:
@@
Here $f$ denotes the ellipsoidal parameterization and $\mathcal{F}$ is the set of all triangular faces. The normalization factors in the numerator and denominator ensure that a perfectly area-preserving mapping would result in $d_{\text{area}}\equiv 0$. In Fig. 6, we present various genus-0 surface models, the initial ellipsoidal parameterizations, the ellipsoidal parameterizations obtained by our EDEM method, and the initial and final log-area errors. It can be observed that the EDEM method performs very well for genus-0 closed surfaces with complex structures. Specifically, by visually comparing the triangle element distributions in the initial ellipsoidal parameterizations and the EDEM parameterization results, it is easy to see that the EDEM method achieves a more uniform distribution on the prescribed ellipsoid. By comparing the initial and final log-area error histograms, we can also see that the log-area error is significantly reduced by our EDEM method. For a more quantitative analysis, Table 2 provides detailed statistics on the performance of our EDEM algorithm for ellipsoidal area-preserving parameterization. We can see that the EDEM method can efficiently reduce the log-area error by over 95% on average when compared to the initial parameterization, and the bijectivity of the parameterization is well-preserved in all examples. Altogether, the experiments show that our EDEM method is capable of computing bijective ellipsoidal area-preserving parameterizations for a large variety of genus-0 closed surfaces.
Figure 6: Ellipsoidal area-preserving parameterization of genus-0 closed surfaces obtained by the EDEM method. Each row shows one example. (a) The David model with elliptic radii $(a,b,c)=(1,1.2,1.4)$. (b) The Buddha model with elliptic radii $(a,b,c)=(1,1,1.4)$. (c) The hippocampus model with elliptic radii $(a,b,c)=(1,1,2.2)$. (d) The twisted ball model with elliptic radii $(a,b,c)=(1,0.8,1.4)$. Left to right: The input surface mesh, the initial ellipsoidal conformal parameterization obtained by the FECM method [30], the EDEM result, the histogram of the log-area error $d_{\text{area}}$ of the initial ellipsoidal parameterization, and the histogram of the log-area error $d_{\text{area}}$ of the final ellipsoidal parameterization.
Table 2: The performance of our EDEM algorithm for ellipsoidal area-preserving parameterization. For each surface, we record the number of triangle elements, the computational time, the elliptic radii $(a,b,c)$ of the target ellipsoidal domain, mean and standard deviation of the initial log-area error $|d_{\text{area}}(f_{0})|$ and the final log-area error $|d_{\text{area}}(f)|$, and the number of overlaps.
Surface # Faces Time (s) $\mathbf{(a,b,c)}$ $|d_{\text{area}}(f_{0})|$ $|d_{\text{area}}(f)|$ # Overlaps
@@
Table 3: The results of the Hippocampus model obtained by our EDEM algorithm with different elliptic radii. Here, for different choices of the elliptic radii $(a,b,c)$ of the target ellipsoid, we record the mean and standard deviation of the final log-area error $|d_{\text{area}}(f)|$, the mean value of the norm of the Beltrami coefficient $|\mu|$, and the number of overlaps.
[…]
@@
It is natural to ask whether the performance of our EDEM method is affected by the shape of the target ellipsoid. Here, we map the Hippocampus model to various target ellipsoids with different elliptic radii using our EDEM method and analyze the area and angle distortions of the resulting mappings. In Table 3, we report the mean and variance of the log-area error, as well as the mean value of the norm of the Beltrami coefficient $|\mu|$. It can be observed that for all combinations of the elliptic radii, all mapping results are folding-free and possess low log-area error. This shows that our EDEM method is capable of computing area-preserving parameterizations of genus-0 closed surfaces onto different prescribed ellipsoids. However, if we also take the angle distortion into consideration, then some differences can be observed. For instance, if we increase the value of $c$, we can see that the angle distortion is further reduced. This experiment suggests that the overall ellipsoidal shape plays an important role and motivates the need for the proposed EDEQ method, which optimizes the ellipsoidal geometry and achieves ellipsoidal density-equalizing quasi-conformal maps.
Figure 7: Ellipsoidal density-equalizing quasi-conformal maps of ellipsoidal surfaces. Each row shows one example. (a) An example with discontinuous input density. Here, the initial elliptic radii are $(a_{0},b_{0},c_{0})=(1,1,2)$ and the final elliptic radii are $(a,b,c)=(1,1.8973,2.8163)$. (b) An example with continuous input density. Here, the initial elliptic radii are $(a_{0},b_{0},c_{0})=(1,2,4)$ and the final elliptic radii are $(a,b,c)=(1,1.9888,4.3110)$. Left to right: The initial ellipsoidal surface color-coded with the initial density, the final EDEQ result color-coded with the initial density, the histogram of the initial density, and the histogram of the final density.
@@ 4.2 Ellipsoidal density-equalizing quasi-conformal map
Next, we test our EDEQ algorithm for computing the ellipsoidal density-equalizing quasi-conformal mappings. We set $\alpha=5$ in the following experiments. In Fig. 7(a), we define different populations at different regions on the ellipsoid to obtain a discontinuous initial density. Note that the maximum density is three times larger than the minimum one. It can be observed in the EDEQ mapping result that the shape of the ellipsoid is changed under the mapping process, while the higher-density domain expands and the lower-density domain contracts. Comparing the initial and final density histograms, it can be observed that the density is highly equalized. Fig. 7(b) shows another example with a continuous initial density defined on an ellipsoid. Under the EDEQ algorithm, the ellipsoidal shape is stretched along the z-direction and becomes more elongated, and different regions are expanded or shrunk according to the given density. From the density histograms, it can again be observed that the mapping result is highly density-equalizing.
@@
Figure 8: Ellipsoidal density-equalizing quasi-conformal maps of ellipsoidal surfaces. For each example, the left figure shows the input surface color-coded with the initial density, and the right figure shows the EDEQ result color-coded with the initial density. (a) An ellipsoid with non-uniformly distributed mesh elements and a prescribed discontinuous density. Here, the initial elliptic radii are $(a_{0},b_{0},c_{0})=(1,1,1.5)$ and the final elliptic radii are $(a,b,c)=(1,1,1.4250)$. (b) An elongated ellipsoid with a prescribed continuous density. Here, the initial elliptic radii are $(a_{0},b_{0},c_{0})=(1,0.6,2)$ and the final elliptic radii are $(a,b,c)=(1,0.9821,2.5162)$. (c) An ellipsoid with non-uniformly distributed mesh elements and a complex prescribed density. Here, the initial elliptic radii are $(a_{0},b_{0},c_{0})=(1,1,1.5)$ and the final elliptic radii are $(a,b,c)=(1,1.0310,1.0783)$. (d) A sphere with a prescribed continuous density. Here, the initial elliptic radii are $(a_{0},b_{0},c_{0})=(1,1,1)$ and the final elliptic radii are $(a,b,c)=(1,1.3095,0.9100)$.
@@
In Fig. 8, we further apply the EDEQ method to the same set of ellipsoidal examples in Fig. 5. It can be observed that our EDEQ method produces ellipsoidal density-equalizing mapping results with both the ellipsoidal geometry and the vertex positions optimized. For a more quantitative analysis, we record the initial and final elliptic radii, the variance of the initial and final densities, the mean value of the Beltrami coefficient $\mu$, and the number of overlaps for all the above-mentioned examples in Table 4. From the table, we can see that the ellipsoidal shapes are effectively optimized under the EDEQ method. Furthermore, in addition to possessing a highly density-equalizing effect as reflected in the variance of the final density, the mapping results also achieve a low conformal distortion. Specifically, note that the average values of the norm of the Beltrami coefficient $\text{mean}(|\mu|)$ in all examples are around 0.2 only, while for general bijective quasi-conformal maps, the range of $|\mu|$ is from 0 to 1, with $|\mu|=0$ indicating that the mapping is conformal and $|\mu|=1$ indicating that there is a singularity. Moreover, the $|\mu|$ value of our method is significantly smaller than that of fixed shape methods, indicating that our method can better preserve the local geometric structure. Hence, we can see that the EDEQ method not only achieves density equalization but also reduces the conformal distortion of the ellipsoidal mappings.
@@
Next, we consider using the EDEQ method to compute the ellipsoidal area-preserving parameterizations for genus-0 closed surfaces. For each surface in Fig. 9, we first conformally map it onto an ellipsoid with some arbitrary elliptic radii using the FECM method [30]. Then, we apply the EDEQ algorithm with the population set as the original face area to obtain the ellipsoidal area-preserving parameterizations. It can be observed that the ellipsoidal shapes are changed significantly under the EDEQ method. Comparing the initial and final log-area error histograms, we can see that the mapping results are highly area-preserving.
Figure 9: Ellipsoidal area-preserving quasi-conformal parameterization of genus-0 closed surfaces. Each row shows one example. (a) The Duck model with initial elliptic radii $(a_{0},b_{0},c_{0})=(1,1,1.2)$ and final elliptic radii $(a,b,c)=(1,1.8797,1.9511)$. (b) The Hippocampus model with initial elliptic radii $(a_{0},b_{0},c_{0})=(1,1,1.5)$ and final elliptic radii $(a,b,c)=(1,0.4487,2.0513)$. (c) The Lion-Vase model with initial elliptic radii $(a_{0},b_{0},c_{0})=(1,1,1.4)$ and final elliptic radii $(a,b,c)=(1,0.47557,1.6682)$. Left to right: The input surface mesh, the initial ellipsoidal conformal parameterization, the final EDEQ result, the histogram of the log-area error $d_{\text{area}}$ of the initial ellipsoidal parameterization, and the histogram of the log-area error $d_{\text{area}}$ of the final ellipsoidal parameterization.
@@
As the choice of the initial elliptic radii is arbitrary, we can even start with a unit sphere and run the EDEQ method to eventually obtain an optimal ellipsoidal parameterization. In Fig. 10, we consider several genus-0 closed surface models and compute an initial spherical conformal parameterization. Then, we apply the proposed EDEQ algorithm to compute an optimal ellipsoidal area-preserving parameterization. It can be observed that throughout the EDEQ iterative process, both the ellipsoidal shapes and the vertex positions on the ellipsoids are changed gradually. For a more quantitative analysis, Table 5 summarizes the performance of our EDEQ algorithm in terms of the final radii, the initial and final log-area error, the conformal distortion, and the number of overlaps. Comparing the initial and final log-area error, it can be observed that the log-area error is significantly reduced in all examples. The mean of the norm of the Beltrami coefficient $|\mu|$ is also small, indicating that the conformal distortion is low. Besides, the number of overlaps is 0 for all examples, which shows that our mapping results are all folding-free. In Table 6, we further compare the performance of the EDEQ method and the EDEM method in terms of the area and conformal distortions, where both methods use the same initial elliptic radii. It can be observed that the EDEQ method achieves a significantly lower conformal distortion when compared to the EDEM method, while the log-area error remains comparable. Besides, we remark that since the EDEQ mappings are diffeomorphisms, they preserve the topological structure between the input genus-0 closed surface and the final ellipsoidal domain. Therefore, the final results of EDEQ will still be genus-0 closed ellipsoidal surfaces with $a,b,c>0$ and will not become a flat disk. In our experiments, we empirically tested convergence and observed the radii to asymptotically approach non-zero values. Altogether, the above experiments demonstrate the effectiveness of our proposed EDEQ method for ellipsoidal density-equalizing quasi-conformal maps.
Table 5: The performance of our EDEQ algorithm for ellipsoidal area-preserving parameterization. For each surface, we start with an initial spherical conformal parameterization and then apply the EDEQ algorithm. The number of triangle elements, the final elliptic radii $(a,b,c)$, the mean and standard deviation of the initial log-area error $|d_{\text{area}}(f_{0})|$ and the final log-area error $|d_{\text{area}}(f)|$, the mean value of the norm of the Beltrami coefficient $|\mu|$, and the number of overlaps are recorded.
@@
[…]